OpenStudy (anonymous):

Substitute your birthday into the equation Square root of the quantity x minus y + m = d, where y is the last two digits of your birth year, m is the month, and d is the day. If you were born on 7/10/1856 like Nikola Tesla, your equation would be Square root of the quantity x minus 56 + 7 = 10. Solve for x and identify if it is an extraneous solution. my bday is 2/6/1997

OpenStudy (anonymous):

x = x y = 97 m = 2 d = 6 ------------------ \[\sqrt{x}-97+2=6\]

OpenStudy (anonymous):

yea then i got is x-97=16

OpenStudy (anonymous):

then i solved for x=113

OpenStudy (anonymous):

then i pluged it in which i got is squaroot 113-96+2=8

OpenStudy (anonymous):